OpenStudy (anonymous):

88+x=3(x+22) Steps please. explain

OpenStudy (anonymous):

x=11

OpenStudy (anonymous):

88+x=3x+66 (expand using distributive law0 88-66=3x-x (bring like terms together) 22=2x 11=x

OpenStudy (anonymous):

Steps using axioms of the real number field: 1) RHS distributive law \[88 + x = 3x + 66\] 2) Existence of an additive inverse \[88 + x + (-88) = 3x + 66 + (-88)\] 3) Commutative property of addition \[88 + (-88) + x = 3x + 66 + (-88)\]\[x = 3x - 22\] 4) Existence of an additive inverse and commutative property of addition \[x + (-3x) = 3x + (-3x) - 22\]\[x - 3x = - 22\] 5) Distributive property \[(1 - 3)x = -22\]\[-2x = -22\] 6) Existence of a multiplicative inverse and commutative property of multiplication \[-2x\cdot(-2^{-1}) = -22\cdot(-2^{-1})\]\[-2\cdot(-2^{-1})\cdot x = -22\cdot(-2^{-1})\]\[x = 11\]